OpenStudy (anonymous):

Solve for x: |3x + 2| = 14 x = 4 and x = −4 x = 4 and x = -16/3 x = 28/3 and x = -16/3 No Solutions

OpenStudy (anonymous):

=[3x+2]=14 =3x=14-2 =3x=12 =x=4

OpenStudy (jhannybean):

@Amolg That is not correct.

OpenStudy (anonymous):

so how ?

OpenStudy (anonymous):

@Jhannybean can u know me for batter understanding this

OpenStudy (jhannybean):

|3x+2| = 14 by opening up this absolute value, we are considering values of x to the left and right side of this function. 3x +2 = -14 Or 3x +2 = 14 3x = -16 Or 3x = 12 x = ? Or x =?

OpenStudy (anonymous):

i wrote it wrong

OpenStudy (anonymous):

@supanick14 ques is |3x + 2| = 14 not 3x +2 = -14

OpenStudy (jhannybean):

O_o

OpenStudy (jhannybean):

No,you have a correct answer there, I see it. Which one of these would you get after solving for x? x = 4 and x = −4 x = 4 and x = -16/3 x = 28/3 and x = -16/3

OpenStudy (anonymous):

i just wrote the answers wrong

OpenStudy (anonymous):

what is the answer i dont know

OpenStudy (jhannybean):

3x= - 16 Solve for x 3x = 12 Solve for x

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

yes thats right ans i think

OpenStudy (anonymous):

[3x+2]=14 =3x=14-2 =3x=12 =x=12/3 =x=4

OpenStudy (anonymous):

againg i same

OpenStudy (jhannybean):

yu will have TWO answers. Like i mentioned in my above post. 3x= -16 divide both sides by 3 3x/3 =-16/3 Simplify x =-16/3 --------------------------------------------------------------------- 3x = 12 divide both sides by 3 3x/3 = 12/3 x = 12/3 = 4.

OpenStudy (anonymous):

ok lol